a two digit number is 6 times the sum of its digits. the tens digit is 1 more than the units digit. find the number.
Answer by xcentaur(357)

let the digits of the number be x and y
then the number is (10x+y)


Given the two digit number is 6 times the sum of its digits.
Sum of its digits = (x+y)
From given statement we get
(10x+y)=6(x+y)..................[1]


Also we have,the tens digit is 1 more than the units digit.
this means that x=y+1...........[2]


From equations one and two we get,
10x+y=6(x+y)
10x+y=6x+6y
10x-6x=6y-y
4x=5y
x=5y/4...........................[3]


we get x=5y/4=y+1
in other words
5y/4=y+1
5y=4(y+1)
5y=4y+4
5y-4y=4
y=4


y=4
x=y+1=5


Then the number is = (10x+y) = 54


Hope this helps